OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of one half. y - 6 = one half(x + 3) y = one halfx - 6 y + 3 = one half(x - 6) x - 2y = 12

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

for starters, we need to know what point slope form looks like in general. any ideas?

OpenStudy (anonymous):

um not really

OpenStudy (amistre64):

what line form equations do you know of? maybe we can narrow it down some

OpenStudy (anonymous):

give ,me a sec let me see ,my notes

OpenStudy (anonymous):

slope intercept form

OpenStudy (amistre64):

i see a slope intercept form up there, so we know thats not the answer. do you see which one we can get rid of?

OpenStudy (anonymous):

yea D

OpenStudy (amistre64):

D is in standard form, B is slope intercept also, the names of the form tend to define what we use to create them. slope intercept is made using the slope and the y intercept. point slope is made by using any given point, and the slope, to construct it with. we want an option that uses the point values and the slope value

OpenStudy (anonymous):

so A is the answer right?

OpenStudy (amistre64):

yes :)

OpenStudy (anonymous):

thanks!

OpenStudy (amistre64):

notice that a line has the same slope between any 2 points. given a point (x,y) and a point (a,b) we can use the slope formula to construct a line \[\frac{y-b}{x-a}=m\] mathing out the bottom converts it to point slope form \[{y-b}=m({x-a})\] and my computer is sooo slow at the moment

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

help me out wit another one? i tagged u already

OpenStudy (amistre64):

i havent got the notif of the tag yet. but ill bee there soon. others have tagged me as well.